• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 00:12:03 on March14,2006

    Entry number 166563

    keyword=swing shift summary

    Shift Leader: A. Puckett
    Target Operator: J. Miller

    16:06 - Took over from day shift. Production running is going smoothly. Current run # is 2937. At 18:25 we took away beam for an NMR polarization measurement. Result is 39.8%. At 18:37 we took the beam back at 7 muA, resumed production running. Hall B asks to do IHWP change, expect this to take place at about 23:45.

    Half-wave plate change takes place at 23:45. Bogdan wants to do HV change, down for a few minutes after IHWP change

    Production Runs
    Run # Start Stop Nevents Replay OK?
    2937 15:37 16:59 3M Y
    2938 17:00 18:24 3M Y
    2939 18:35 19:57 3M Y
    2940 19:57 20:51 2M Y
    2941 20:51 21:49 2M Y
    2942 21:50 22:44 2M Y
    2943 22:44 23:39 2M Y
    2944 23:56

    Compton Runs: (Start/Stop times are in Compton time, not necessarily the same as HALOG time)
    Run # Start Stop S/N Pe*Pg Replay OK?
    13084 13:56 16:03 12.5 82.6 Y
    13085 16:07 18:15 22.5 83.6 Y
    13086 18:18 20:25 21.7 88.7 Y
    13087 20:34 22:40 26.4 82.8 Y
    13088 22:48 23:43 21.6 76.0 Last few minutes were after 1/2 wave plate change, re-analyze by expert?
    13089 00:03

    Parity Runs:
    Run # Start Stop N Aq
    7074 16:32 18:25 200K 1.679
    7075 18:37 23:19 500K -1.772
    7076 23:52